OpenStudy (anonymous):

(-6)(-2)(-1) = a0 please help

OpenStudy (anonymous):

6*2*1 = 12 (-6)*2*1 = -12 (-6)*(-2)*1 = 12 (-6)*(-2)*(-1) = -12 To each minus sign you change the sign of the answers one time. Do you get it?

OpenStudy (anonymous):

sorry i do not

OpenStudy (anonymous):

When you are multipling by -1 you change the number sign. Example: (-1)*3 = -3. So you can try to look your problem on that way: (-6)*(-2)*(-1) = [(-1)*6]*(-2)*(-1) = [(-1)*6]*[(-1)*2]*(-1) = (-1)*(-1)*(-1)*6*2 = (-1)*(-1)*[(-1)*6*2] = (-1)*(-1)*[(-1)*12] = (-1)*(-1)*(-12) = (-1)*[(-1)*(-12)] = (-1)*(12) = -12